Sarah leaves Seattle for New York in her car, averaging 80 mph across open country. One hour later a plane leaves Seattle for new york following the same route and flying 400 mph. how long will it be before the plane overtakes the car?

Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)

Sarah leaves Seattle for New York in her car, averaging 80 mph across open country. One hour later a plane leaves Seattle for new york following the same route and flying 400 mph. how long will it be before the plane overtakes the car?
----------------
Sarah is 80 miles ahead (80 * 1)
The plane gains on her at 320 mph (400 - 80)
80/320 = 1/4 hour
-------------------

Answer by ewatrrr(24785)


Hi
Let x represent the amount of time before the plane overtakes the car
D = r*t, distance traveled the same for both car and plane
Question States***
80mph(x + 1hr) = 400mph*x
solving for x
80 = 320x
x = 1/4hr, the amount of time before the plane overtakes the car
CHECKING our Answer
80*5/4 = 1/4*400 = 100mi (the distance traveled by both when plane passes)